A cylindrical silo has a height of 15.5 meters and a diameter of 4.2 meters if 1 can of paint covers 10 square meters how many cans must be bought to paint the curved face with one coat


1
Expert's answer
2022-04-08T05:35:30-0400

Area of the curved face: "A=\pi dh=\pi*4.2*15.5=204.5 m^2."

Number of cans: "N=\frac{A}{10}=\frac{204.5}{10}=20.45\;cans\approx 21 \;cans."